A buffer solution is composed of 0.380 M HA, a weak monoprotic acid, and 0.760 M NaA, the sodium salt of the acid. The solution has a pH of 4.10. What is the Ka of the weak acid, HA?

tks!!!!!!!!!!!!!
Answer #1

use the Handerson equation find the pKa first from that find the Ka value

pH = pKa + log(NaA/HA)

pH is given = 4.10

NaA = 0.76 M,, HA = 0.38

put all these in the above equation

4.10 = pKa + log(0.76/0.38)

pKa = 4.10-0.3010

pKa = 3.8

now use the formula

ka = 10-pKa

Ka = 10-3.8

Ka = 1.58 x 10-4